Determine the maximum intensity w such that the beam does not pivot downward more than 0.050 5M 375 N/m 10 com Sm 7m
Answer #1

m 175 Nim Kubomok som kan 7m Given that maximum pivol is 0.05°. Let us consider that beam has pivoled then the figure well lodue to deflection Stiffness of spring K= 75 Nilm upward force applied by Spring F = kxx = 75% 5.235x10-3 N. F = 0.3926 N. F aBoth for the moments deviation to 1 exceed and 2 must be equal 0.05 - MA = MAI 2.3556 = 101-5 mi .. lw= 0·0232 Nlm 1 - Answer